OpenStudy (anonymous):

what multiplies to be -4 and adds to be -4

OpenStudy (anonymous):

let x and y be the two numbers where xy = -4 and x+y = -4 x = -4 -y y(-4-y) = -4 -4y-y^2 = -4 y^2 +4y -4 = 0 Solve for y then for x